Cycle Center has bicycles and tricycles in the storeroom. There are at least three of each, and there are more bicycles than tricycles. There are 28 wheels altogether. How many bicycles and tricycles are in the storeroom?

Answer by Edwin McCravy(20054)

Cycle Center has bicycles and tricycles in the storeroom. There are at least
three of each, and there are more bicycles than tricycles. There are 28
wheels altogether. How many bicycles and tricycles are in the storeroom?
b = number of bicycles
t = number of tricycles
There are 28 wheels altogether.
2b + 3t = 28

2 is the smaller coefficient in absolute value. So write each number 
which is not a multiple of 2 in terms of its nearest multiple of 2.

So we only need to write the 3 as 2+1

2b + (2+1)t = 28
2b + 2t + t = 28

Divide each term by 2

 b + t + t/2 = 14

Isolate any fractions

         t/2 = 14 - b - t

The right side is an integer, so the left side must be equal to the
same integer.  Let that integer be K

t/2 = K,    14 - b - t = K

  t = 2K,   14 - b - 2K = K

                14 - 3K = b
There are more bicycles than tricycles:
      b > t
14 - 3K > 2K
     14 > 5K
    2.8 > K
      K ≤ 2 since K is an integer
There are at least three of each
 
t = 2K ≥ 3,   b = 14 - 3K ≥ 3
     K ≥ 1.5,         -3K ≥ -14
     K ≥ 2              K ≤ 14/3
                        K ≤ 4.333...
                        K ≤ 4
(since K is an integer)

Since K ≤ 2 and K ≥ 2, then K = 2

Thus t = 2K = 2(2) = 4 tricycles and
     b = 14 - 3K = 14 - 3(2) = 14-6 = 8 bicycles.

Edwin